OpenStudy (anonymous):

@Vijeya3 Last One ;) Factor. x^2 + 25x - 100 A. (x - 2)(x + 50) B. (x - 5)(x + 20) C. (x + 4)(x - 25) D. Prime

OpenStudy (mathmale):

You could check each set of factors by multiplying them together (FOIL method). For example: (x-5)(x+20)= x^2 + 20x - 5x - 100 = x^2 + 15x - 100. Is this the same as the original quadratic?